# Understanding the Application of Utility Theory in Robotics and Artificial Intelligence: A Survey

### v1

@@ IV-B Cooperative Decision-making and Learning
Furthermore, existing cooperative decision-making models use Markov decision process (MDP) and its variants [86], game-theoretic methods, and swarm intelligence [84]. They mostly involve using Reinforcement Learning (RL) and Recurrent Neural Networks (RNN) to find optimal or suboptimal action sequences based on current and previous states for achieving independent or transferred learning of decision-making policies [87, 88].
Considering decision-making in the context of cooperative MAS, the learning process can be centralized or decentralized. [89] divides it into two categories: team learning and concurrent learning.
@@
Figure 8: End-to-End learning of coordinated 2 vs 2 humanoid football in MuJoCo multi-agent soccer environment
@@
In contrast, if we do not divide the team rewards equally, evaluating each agent’s performance is based on individual behaviors, which means agents do not have the motivation to cooperate and tend to greedy behaviors; these methods are the so-called local reward. Through studies different credit assignment policies, [92, 93] argue that local reward leads to faster learning rates, but not necessarily to better results than global reward. Because local rewards increase the homogeneity of the learning group, he suggests that credit assignment selection relies on the desired degree of specialization. Furthermore, [94] argues that individual learning processes can be improved by combining separate local reinforcement with types of social reinforcement.
@@
The dynamics of learning consider the impact of co-adaptation on the learning processes. Assuming agents work in dynamic changing and unstructured environments, they need to constantly track the shifting optimal behavior or strategy adapting to various situations, especially considering the agents may change other group members’ learning environments. Evolutionary game theory provides a common framework for cooperative MAS learning. [95, 96] studied the properties of cooperative coevolution and [97] visualize basins of attraction to Nash equilibria for cooperative coevolution.
@@
Moreover, lots of research in concurrent learning involve game theory to investigate MAS problems [98, 8]. Especially for Nash equilibria, it provides a joint strategy for each group member, which lets individuals not have the motivation to shift their strategy or action in the current situation in terms of a better reward. In the full-cooperative scenario with global reward, the reward affects the benefits of each agent and only needs to consider the globally optimal Nash equilibrium. However, the relationships among agents are not clear for more general situations, which combine the cooperative and non-cooperative scenarios. In other words, individual rewards do not directly relate to the team reward, so the general-sum game [8] is the cooperative learning paradigm.
IV-C From Individual Utility to Social Welfare
@@
As we discussed above, the reward same as the utility, reflects the agents’ diverse intrinsic motivations and needs. Individual and team rewards also directly describe the relationships between each group member and the team. We are taking MAS football and the Robot-soccer domain as an example. When considering multi-agent systems (MAS) cooperation as a team, an individual agent must first master essential skills to satisfy low-level needs (safety, basic, and capability needs) [51]. Then, it will develop effective strategies fitting middle-level needs (like teaming and collaboration) to guarantee the systems’ utilities [52]. Through learning from interaction, MAS can optimize group behaviors and present complex strategies adapting to various scenarios and achieving the highest-level needs, and fulfilling evolution. By cooperating to achieve a specific task, gaining expected utility (reward), or against the adversaries decreasing the threat, intelligent agents can benefit the entire group development or utilities and guarantee individual needs and interests. It is worth mentioning that [99] developed the end-to-end learning implemented in the MuJoCo multi-agent soccer environment [100], which combines low-level imitation learning, mid-level, and high-level reinforcement learning, using transferable representations of behavior for decision-making at different levels of abstraction. Figure 8 represents the training process.
@@
However, suppose we extend the scalar to the entire system, like community and society, and consider optimizing the overall social welfare across all agents. In that case, we have to introduce the social choice utility. In this setting, agents have different agendas driven by their needs, leading to complex dynamic behaviors when interacting, which causes the utility hard to predict, let alone optimize. To solve this problem, we usually assume that agents are self-interested and optimize their utilities from the perspective of socially favorable desires and needs [58]. By defining the social choice function representing social welfare reflecting each agent’s needs and desired outcome, we can design a system of payments making the joint policy converge to the desired outcome [56].
For example, the government controls the parameters of tenders to balance the different requirements of projects for social welfare. Also, [101] discussed computing a convex coverage set for a cooperative multi-objective multi-agent MDP balancing traffic delays and costs as a posteriori in traffic network maintenance planning.
@@
From another perspective, considering individual rewards and utilities from the standpoint of social choice, they are weighed and summed up, consisting of a social welfare function. In other words, the social welfare function relies on individual values or return rewards and its utility computing model or mechanism. For example, [102] discussed which properties must satisfy a multi-criteria function so it can be used to determine the winner of a multi-attribute auction. Here, social welfare may depend on attributes of the winning bids, as well as a fair outcome in terms of payments to the individual agents, that together with the costs the agents need to support to execute their bids, if chosen, typically determine the individual utilities [58].
@@
Generally speaking, the design of the social welfare function aims to find a mechanism for making agents’ utility function transparently reflect their innate values and needs, which can optimize the joint policy concerning sustainable social welfare and guarantee individual development.
V Trust among Agents
@@
In Automation, authors in [106] describe trust in human-automation relationships as the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability. The trustor here is a human and the trustee usually is the automation system or an intelligent agent like robot. Those systems’ primary purpose is to assess and calibrate the trustors’ trust beliefs reflecting the automation system’s ability to achieve a specific task.
In Computing and Networking, the concept of trust involves many research fields, such as artificial intelligence (AI), human-machine interaction, and communication networks [107, 108, 109]. They regard an agent’s trust as a subjective belief which represents the reliability of the other agents’ behaviors in a particular situation with potential risks. Their decisions are based on learning from the experience to maximize its interest (or utility) and minimize risk [110]. Especially in the social IoT domain [111, 112], trust is measured based on the information accuracy and agents’ intentions (friendly or malicious) according to the current environment to avoid attacks on the system. [113] proposes a decision-making model named MATE, which applies the expected-utility decision rule to decide whether or not pervasive devices are in their node interests to load a particular component.
In MAS, trust is defined by mainly two major ways. One way is reliable relations between human operators and the MAS/MRS. In this context, trust reflects human expectations and team performance assessments [114, 115]. The primary research includes using the trust to measure the satisfactory degree of robot team performance that hybrid considering in-process behaviors and final mission accomplishments. In this process, the visible and invisible behaviors regarding humans supervising robots and robots’ behaviors will be investigated to decide the most influential factors. Especially investigating human trust in robots will help the robots to make decisions or react to emergencies, such as obstacles, when human supervision is unavailable [116, 117].
@@
It is worth noticing that reputation and trust mechanisms have gradually become critical elements in MAS design since agents depend on them to evaluate the behavior of potential partners in open environments [120]. As an essential component of reputation, trust generates the reputation from agents’ previous behaviors and experiences, which becomes an implicit social control artifact [121] when agents select partners or exclude them due to social rejection (bad reputation).
@@
Moreover, with the development of the explanation of machine learning, such as explainable AI, although explanation and trust have been intertwined for historical reasons, the utility of machine learning explanations will have more general and consequential [127]. Especially the utility of an explanation must be tied to its appropriate trust model.
Although modeling trust has been studied from various perspectives and involves many different factors, it is still challenging to develop a conclusive trust model that incorporates all of these factors. Therefore, future research into trust modeling needs to be more focused on developing general trust models based on measures other than the countless factors affecting trust [122]. Specifically, few works from the literature considered trust in socio-intelligent systems from an AI agent perspective to model trust with agents having similar needs and intrinsic values.
Recently, [128] proposed a novel trust model termed relative needs entropy (RNE) by combining utility theory and the concept of relative entropy based on the agent’s needs hierarchy [51, 52] and innate values. Similar to information entropy, it defines the entropy of needs as the difference or distance of needs distribution between agents in a specific scenario for an individual or group.
From a statistical perspective, the RNE can be regarded as calculating the similarity of high-dimensional samples from the agent needs vector. A lower RNE value means that the trust level between agents or groups is higher because their needs are well-aligned and there is low difference (distance) in their needs distributions. Similarly, a higher RNE value will mean that the needs distributions are diverse, and there exists a low trust level between the agent or groups because of their misalignment in their motivations.
@@
VI Human-Robot Interaction
As the higher-level intelligent creature, humans represent more complex and diversified needs such as personal security, health, friendship, love, respect, recognition, and so forth. In this regard, humans are highly flexible resources due to the high degree of motion of the physical structure and the high level of intelligence and recognition [129]. When we consider humans and robots work as a team, organizing their needs and getting a common ground is a precondition for human-robot collaboration in complex missions and dynamic changing environments. Considering building more stable and reliable relationships between humans and robots, [130] depicts an updated model in HRI termed Human Robot Team Trust (HRTT). It considers the human as a resource with specific abilities that the resource can gain or lose trust with training and realizes the robots as resources with known performance attributes included as needs in the design phase of the collaborative process.
Due to the constantly increasing need for robots to interact, collaborate and assist humans, Human-Robot Interaction (HRI) poses new challenges, such as safety, autonomy, and acceptance issues [131]. From a robot needs perspective [51], it first needs to guarantee human security and health, such as avoiding collision with humans, protecting them from risks, etc. Moreover, in the higher level teaming needs, robots should consider human team members’ specialty and capability to form corresponding heterogeneous Human-Robot team adapting specific missions automatically [52]. Furthermore, efficient and reliable assistance is essential to the entire mission process in HRI. More importantly, designing an Interruption Mechanism can help humans interrupt robots’ current actions and re-organize them to fulfill specific emergency tasks or execute some crucial operations manually.
@@ VI Human-Robot Interaction
For Multi-agent HRI systems, they require more complex control strategies to coordinate several agents that may be dissimilar to one another (in roles, communication modes etc.), can involve interactions that connect more than two agents at once, and need special attention to address any conflicts that may arise from their interactions [138]. Specifically, the human would give high-level commands to the robot group or control them individually using a screen-based interface, and robots need to adapt to the operator’s actions or optimize some utility function (computational characteristics like task requirements, teaming needs, etc.) (See Fig. 2).
Moreover, building the shared mental models (SMM) enable team members to draw on their own well-structured common knowledge to select actions that are consistent and coordinated with those of their teammates, which are strongly correlated to team performance [139]. [140] designs the human-robot cross-training inspired by SMM to compute the robot policy aligned with human preference (needs) by iteratively switching roles between humans and robots to learn a shared plan for a collaborative task. Furthermore, considering humans infer the robot’s capabilities and partially adapt to the robot, [141] presents a game-theoretic model of human partial adaptation to the robot, where the human responds to the robot’s actions by maximizing a reward function that changes stochastically over time, capturing the evolution of their expectations of the robot’s capabilities.
@@
VII Insights and Challenges
Currently, with the tremendous growth in AI technology, Robotics, IoT, and high-speed wireless sensor networks (like 5G), it gradually forms an artificial ecosystem termed artificial social systems that involves entire AI agents from software entities to hardware devices [53]. How to integrate artificial social systems into human society and coexist harmoniously is a critical issue for the sustainable development of human beings, such as assistive and healthcare robotics [145], social path planning and navigation [146], search and rescue [147, 52], and autonomous driving [148]. Especially the future factory is likely to utilize robots for a much broader range of tasks in a much more collaborative manner with humans, which intrinsically requires operation in proximity to humans, raising safety and efficiency issues [139].
At this point, building a robust, stable, and reliable trust network between humans and AI agents to evaluate agents’ performance and status in a common ground is the pre-condition for efficient and safe interaction in their collaboration. Here, utility is the key concept representing the individual agent’s innate values, needs and motivations over states of the world and is described as expected rewards. It measures sensitivity to the impact of actions on trust and long-term cooperation and is efficient enough to allow AI agents like robots to make real-time decisions [149]. For example, self-driving cars might be the first widespread case of trustworthy robots designed to earn trust by demonstrating how well they follow social norms.
From the learning perspective, the individual learning model can be regarded as constructing models of the other agents, which takes some portion of the observed interaction history as input, and returns a prediction of some property of interest regarding the modeled agent [150]. Taking human society as an example, human beings represent more complex and diversified needs, such as personal security, health, friendship, love, respect, recognition, and self-actualization. These diverse needs can be expressed as corresponding rewards or utility mechanisms driving people to adopt various strategies and skills to achieve their requirements. Especially the multi-layer and diverse needs of people are the foundation for describing their complex and dynamic relationships, such as trust.

### v3

@@ 4.2 Cooperative Decision-making and Learning
Furthermore, existing cooperative decision-making models use the Markov decision process (MDP) and its variants [292], game-theoretic methods, and swarm intelligence [216]. They mostly involve using Reinforcement Learning (RL) and Recurrent Neural Networks (RNN) to find optimal or suboptimal action sequences based on current and previous states for achieving independent or transferred learning of decision-making policies [295, 214]. Considering decision-making in the context of cooperative MAS, the learning process can be centralized or decentralized. [200] divides it into two categories: team learning and concurrent learning.
@@
In contrast, if we do not divide the team rewards equally, evaluating each agent’s performance is based on individual behaviors, which means agents do not have the motivation to cooperate and tend to greedy behaviors; these methods are the so-called local reward. Through studies of different credit assignment policies, [20, 18] argue that local reward leads to faster learning rates, but not necessarily to better results than global reward. Because local rewards increase the homogeneity of the learning group, he suggests that credit assignment selection relies on the desired degree of specialization. Furthermore, [165] argues that individual learning processes can be improved by combining separate local reinforcement with types of social reinforcement.
Figure 22: The utility rewards classification in multi-objective multi-agent decision-making with different setting [213].
@@
The dynamics of learning consider the impact of co-adaptation on the learning processes. Assuming agents work in dynamic changing and unstructured environments, they need to constantly track the shifting optimal behavior or strategy adapting to various situations, especially considering the agents may change other group members’ learning environments. [76] proposes a method using probabilistic and learning methods for social robots’ decision-making based on affective generation. Evolutionary game theory provides a common framework for cooperative MAS learning. [79, 271] studied the properties of cooperative coevolution and [201] visualize basins of attraction to Nash equilibria for cooperative coevolution. More recently, [280] introduces a general MAS innate-values reinforcement learning architecture from the perspective of individual preferences.
@@
Moreover, lots of research in concurrent learning involve game theory to investigate MAS problems [91, 234]. Especially for Nash equilibria, it provides a joint strategy for each group member, which prevents individuals from being motivated to shift their strategy or action in the current situation in terms of a better reward. In the full-cooperative scenario with global reward, the reward affects the benefits of each agent and only needs to consider the globally optimal Nash equilibrium. However, the agents’ relationships are unclear for more general situations, which combine the cooperative and non-cooperative scenarios. In other words, individual rewards do not directly relate to the team reward, so the general-sum game [234] is the cooperative learning paradigm.
4.3 From Individual Utility to Social Welfare
@@
As we discussed above, the reward same as the utility, reflects the agents’ diverse intrinsic motivations and needs. Individual and team rewards also directly describe the relationships between each group member and the team. We are taking MAS football and the Robot-soccer domain as an example. When considering multi-agent systems (MAS) cooperation as a team, an individual agent must first master essential skills to satisfy low-level needs (safety, basic, and capability needs). Then, it will develop effective strategies fitting middle-level needs (like teaming and collaboration) to guarantee the systems’ utilities. Through learning from interaction, MAS can optimize group behaviors and present complex strategies adapting to various scenarios and achieving the highest-level needs, and fulfilling evolution. By cooperating to achieve a specific task, gaining expected utility (reward), or against the adversaries decreasing the threat, intelligent agents can benefit the entire group development or utilities and guarantee individual needs and interests. [148] developed the end-to-end learning implemented in the MuJoCo multi-agent soccer environment, which combines low-level imitation learning, mid-level, and high-level reinforcement learning, using transferable representations of behavior for decision-making at different levels of abstraction. Figure 25 represents the training process.
@@
However, suppose we extend the scalar to the entire system, like community and society, and consider optimizing the overall social welfare across all agents. In that case, we have to introduce the social choice utility. In this setting, agents have different agendas driven by their needs, leading to complex dynamic behaviors when interacting, which causes the utility hard to predict, let alone optimize. To solve this problem, we usually assume that agents are self-interested and optimize their utilities from the perspective of socially favorable desires and needs [213]. By defining the social choice function representing social welfare reflecting each agent’s needs and desired outcome, we can design a system of payments making the joint policy converge to the desired outcome [258]. Moreover, [229] discussed computing a convex coverage set for a cooperative multi-objective multi-agent MDP balancing traffic delays and costs as a posteriori in traffic network maintenance planning.
@@
From another perspective, considering individual rewards and utilities from the standpoint of social choice, they are weighed and summed up, consisting of a social welfare function. In other words, the social welfare function relies on individual values or return rewards and its utility computing model or mechanism. [208] discussed which properties must satisfy a multi-criteria function so it can be used to determine the winner of a multi-attribute auction. Here, social welfare may depend on attributes of the winning bids, as well as a fair outcome in terms of payments to the individual agents, that together with the costs the agents need to support to execute their bids, if chosen, typically determine the individual utilities [213].
@@
Generally speaking, the design of the social welfare function in MAS aims to find a mechanism for making agents’ utility function transparently reflect their innate values and needs, which can optimize the joint policy concerning sustainable social welfare and guarantee individual development.
5 Trust among Agents
@@
In Automation, authors in [138] describe trust in human-automation relationships as the attitude that an agent will help achieve an individual’s goals in a situation characterized by uncertainty and vulnerability. The trustor here is a human and the trustee usually is the automation system or an intelligent agent like robots. Those systems’ primary purpose is to assess and calibrate the trustors’ trust beliefs reflecting the automation system’s ability to achieve a specific task.
In Computing and Networking, the concept of trust involves many research fields, such as artificial intelligence (AI), human-machine interaction, and communication networks [59, 233, 263]. They regard an agent’s trust as a subjective belief that represents the reliability of the other agents’ behaviors in a particular situation with potential risks. Their decisions are based on learning from the experience to maximize its interest (or utility) and minimize risk [58]. Especially in the social IoT domain [56, 178], trust is measured based on the information accuracy and agents’ intentions (friendly or malicious) according to the current environment to avoid attacks on the system. [212] proposes a decision-making model named MATE, which applies the expected-utility decision rule to decide whether or not pervasive devices are in their node interests to load a particular component.
Figure 23: Illustration of The MAS (UAV) Working in Adversarial Environments with Several Typical Scenarios
In MAS, trust is defined by mainly two major ways. One way is reliable relations between human operators and the MAS/MRS. In this context, trust reflects human expectations and team performance assessments. The primary research includes using the trust to measure the satisfactory degree of robot team performance that hybrid considering in-process behaviors and final mission accomplishments. In this process, the visible and invisible behaviors regarding humans supervising robots and robots’ behaviors will be investigated to decide the most influential factors. Especially investigating human trust in robots will help the robots to make decisions or react to emergencies, such as obstacles, when human supervision is unavailable.
@@
It is worth noticing that reputation and trust mechanisms have gradually become critical elements in MAS design since agents depend on them to evaluate the behavior of potential partners in open environments [206]. As an essential component of reputation, trust generates the reputation from agents’ previous behaviors and experiences, which becomes an implicit social control artifact [62] when agents select partners or exclude them due to social rejection (bad reputation). Fig. 26 illustrates the classification of trust modeling approaches from individual to multi-agent systems from different perspectives.
@@
Moreover, with the development of the explanation of machine learning, such as explainable AI, although explanation and trust have been intertwined for historical reasons, the utility of machine learning explanations will have more general and consequential [68]. Especially the utility of an explanation must be tied to its appropriate trust model. Although modeling trust has been studied from various perspectives and involves many different factors, it is still challenging to develop a conclusive trust model that incorporates all of these factors. Therefore, future research into trust modeling needs to be more focused on developing general trust models based on measures other than the countless factors affecting trust [121]. Specifically, few works from the literature considered trust in socio-intelligent systems from an AI agent perspective to model trust with agents having similar needs and intrinsic values.
Recently, [284] proposed a novel trust model termed relative needs entropy (RNE) by combining utility theory and the concept of relative entropy based on the agent’s needs hierarchy [282, 283] and innate values. Similar to information entropy, it defines the entropy of needs as the difference or distance of needs distribution between agents in a specific scenario for an individual or group Fig. 27. From a statistical perspective, the RNE can be regarded as calculating the similarity of high-dimensional samples from the agent needs vector. A lower RNE value means that the trust level between agents or groups is higher because their needs are well-aligned and there is a low difference (distance) in their needs distributions. Similarly, a higher RNE value will mean that the needs distributions are diverse, and there exists a low trust level between the agent or groups because of the misalignment in their motivations.
For future cognitive robotics research, AI robots learn and adapt to human needs and maintain trust and rapport among robots, which are critical for task efficiency and safety improvement [193]. And we list potential directions as follow:
  • Combining the specific capabilities and needs of robots and humans, calculating sensible strategies to efficiently organize the entire group collaboration and fulfill the corresponding mission.
@@
Figure 24: Illustration of five pursuers catching one evader in Pursuit-Evasion game through the game-theoretic utility tree (GUT) with circle formation (level 1) and PP tactic (level 2) in Robotarium [285].
6 Human-Robot Interaction
As the higher-level intelligent creature, humans represent more complex and diversified needs such as personal security, health, friendship, love, respect, recognition, and so forth. In this regard, humans are highly flexible resources due to the high degree of motion of the physical structure and the high level of intelligence and recognition [255]. When we consider humans and robots working as a team, organizing their needs and getting a common ground is a precondition for human-robot collaboration in complex missions and dynamically changing environments. Considering building more stable and reliable relationships between humans and robots, [224] depicts an updated model in HRI termed Human-Robot Team Trust (HRTT). It considers the human as a resource with specific abilities that the resource can gain or lose trust with training and realizes the robots as resources with known performance attributes included as needs in the design phase of the collaborative process.
Due to the constantly increasing need for robots to interact, collaborate and assist humans, Human-Robot Interaction (HRI) poses new challenges, such as safety, autonomy, and acceptance issues [293]. From a robot needs perspective [282], it first needs to guarantee human security and health, such as avoiding collision with humans, protecting them from risks, etc. Moreover, in the higher level teaming needs, robots should consider human team members’ specialty and capability to form corresponding heterogeneous Human-Robot teams adapting specific missions automatically [283]. Furthermore, efficient and reliable assistance is essential to the entire mission process in HRI. More importantly, designing an Interruption Mechanism can help humans interrupt robots’ current actions and re-organize them to fulfill specific emergency tasks or execute some crucial operations manually.
@@ 6 Human-Robot Interaction
Figure 25: End-to-End learning of coordinated 2 vs 2 humanoid football in MuJoCo multi-agent soccer environment [288]
For Multi-agent HRI systems, they require more complex control strategies to coordinate several agents that may be dissimilar to one another (in roles, communication modes etc.), can involve interactions that connect more than two agents at once, and need special attention to address any conflicts that may arise from their interactions [66]. Specifically, the human would give high-level commands to the robot group or control them individually using a screen-based interface, and robots need to adapt to the operator’s actions or optimize some utility function (computational characteristics like task requirements, teaming needs, etc.).
Moreover, building the shared mental models (SMM) enable team members to draw on their own well-structured common knowledge to select actions that are consistent and coordinated with those of their teammates, which are strongly correlated to team performance [249]. [191] designs the human-robot cross-training inspired by SMM to compute the robot policy aligned with human preference (needs) by iteratively switching roles between humans and robots to learn a shared plan for a collaborative task. Furthermore, considering humans infer the robot’s capabilities and partially adapt to the robot, [190] presents a game-theoretic model of human partial adaptation to the robot, where the human response to the robot’s actions by maximizing a reward function that changes stochastically over time, capturing the evolution of their expectations of the robot’s capabilities. Recently, [278] introduces the architecture of edge cognitive computing by combining human experts and assisted robots collaborating in the same framework to achieve a seamless remote diagnosis, round-the-clock symptom monitoring, emergency warning, therapy alteration, and advanced assistance.
@@
7 Insights and Challenges
Figure 26: A classification of approaches in trust modeling from individual to multi-agent systems
As we discussed above, there has been remarkable progress in robotics through utility theory based cognitive modeling, such as intrinsically motivated reinforcement learning systems, diverse values developed and implemented to motivate exploration and learning, and advanced machine learning approaches, like artificial neural networks, reinforcement learning, and Bayesian learning, to build sophisticated value systems. However, several fundamental questions and broad research gaps remain in cognitive robotics: How to build a value system that enables robots to automatically generate motivations in various situations? How can a robot develop an understanding of the structure of its body related to its value system? How can a robot learn the relationships between itself, physical environments, and other agents that inhabit that environment based on its value system?
7.1 Complex Value Systems
For biological entities, their value systems represent the brain as having reused many neural elements (Fig. 8) that evolved originally for sensory pleasures to also mediate higher pleasures [140, 30], which form our reward utility mechanism for novelty, curiosity, competence, concepts of hunger, thirst, and so on. This mechanism motivates complex behaviours and skills, driving them to improve performance to meet their needs and goals. However, we remain unclear about whether our understanding of motivation and value in natural systems is sufficient to inform the design of complex developmental cognitive systems, especially when modeling them in robots. At this point, we face many grand challenges in cognitive modeling regarding value systems:
  • How to build a mechanism to share the complexity of different types of motivations and values from natural systems to artificial systems, such as between humans and robots?
  • How to organize different types of motivations and values to integrate into a single value system for a developmental artificial system?
  • How to let AI agents explore new values or needs to upgrade their value systems for adapting to dynamic changing environments automatically?
  • How to explicitly transfer a specific goal from the value system?
  • How to deal with agents with heterogeneous value systems in their interaction, such as negotiation, consensus, and collaboration?
7.2 Value Systems based Learning
Figure 27: Illustration of The Robots’ Trust based on Robot Needs Hierarchy [284]
From the learning perspective, the individual learning model can be regarded as constructing models of the other agents, which takes some portion of the observed interaction history as input, and returns a prediction of some property of interest regarding the modeled agent [6]. At this point, value systems for developmental cognitive robots have been embedded across a range of robotic platforms, including wheeled mobile robots, humanoids, insectoids, and animal robots. However, when we consider multi-degrees of freedom in robot learning, it becomes both time-inefficient and memory-intensive, even with current state-of-the-art approaches. Therefore, developing approaches that enable robots to autonomously shift their attention from one body part to another and to combine skills across different body parts to learn complex, whole-body behaviors and interact with complex environments are much more critical. Furthermore, another interesting direction is the study of generating behavioural diversity from past experience [29] to shape robots such as animals [65], which is complementary to developmental cognitive robotics. It provides ways to transfer knowledge from one task to new and unexpected tasks in a creative way.